Two equal positive charges q1 q2 = 4.0 μC are located at (1.20 m, 0) and (-1.20 m, 0), = respectively. What are the magnitude and direction of the total electric force that q1 and q2 exerts on a third charge q3= 7.0 μC located at (0, -0.50 m) ?


1
Expert's answer
2022-03-15T16:49:34-0400

"F=\frac{kq_1q_3}{r_1^2+r_2^2}=0.149~N,"

"\phi=\arctan\frac{r_2}{r_1}=67.4°,"

"F_0=\sqrt{{2F^2+2F^2\cos2\phi}}=0.115~N," down.